# PolypDB: A Curated Multi-Center Dataset for Development of AI Algorithms in Colonoscopy

### v1

@@ III-A2 Implementation details
All the methods are implemented using the PyTorch 1.9 [18] framework, which is processed on an NVIDIA GeForce RTX 3090 system. We have used 80% of the dataset for training, 10% for validation and the remaining 10% for testing. For polyp segmentation, we have first resized all the images into $512\times 512$ pixels. Next, we have used a minimal data augmentation, which includes random rotation, horizontal flipping, vertical flipping, and course dropout. A combination of binary cross-entropy loss and dice loss was selected as the loss function with the Adam optimizer, and a learning rate of $1e^{-4}$ was set. We have trained all the models with the same set of hyperparameters for 200 epochs (empirically set) with a batch size of 12. Early stopping and ReduceLROnPlateau was used to prevent the model from overfitting.
For polyp detection, we employed different hyperparameters tailored to optimize the performance of the detection algorithms. At first, we resized all the images to $640\times 480$ and used a simple data augmentation strategy which include random flipping, random rotation, random blur, mixup, mosaic and cutmix. All YOLO models were trained using a uniform set of hyperparameters, with the AdamW optimizer applied at a learning rate of $1\times 10^{-4}$ and a batch size of $16$.
@@ III-B Results
TABLE III: Quantitative results comparison on the PolypDB dataset. The best scores are shown in bold where whereas the second best score is underline.
@@
Figure 2: Qualitative results for the different modalities in the PolypDB dataset.
TABLE IV: Quantitative results comparison on the polyp datasets. The best scores are shown in bold where as the second best score is highlighted in red.
Method MAP50 MAP50-95 MAP75 P R
@@
YOLOv8 [30] 0.659 0.502 0.559 1 0.318
YOLOv10 [31] 0.534 0.416 0.485 0.84 0.5
YOLOv9 [32] 0.688 0.558 0.638 0.846 0.5
@@
YOLOv10 [31] 0.887 0.752 0.875 1 0.853
YOLOv9 [32] 0.856 0.711 0.737 0.937 0.75
YOLOv7 [33] 0.734 0.642 0.734 0.856 0.75
YOLOv6 [34] 0.761 0.608 0.658
@@
YOLOv8 [30] 0.833 0.771 0.833 1 0.667
YOLOv10 [31] 0.995 0.831 0.995 1 0.854
YOLOv9 [32] 0.972 0.878 0.972 0.857 1
@@
YOLOv8 [30] 0.659 0.502 0.559 1 0.318
YOLOv10 [31] 0.534 0.416 0.485 0.84 0.5
YOLOv9 [32] 0.688 0.558 0.638 0.846 0.5
@@
Figure 3: Qualitative results for the different modalities in the PolypDB dataset.
To evaluate the segmentation performance of the dataset, we employed several established segmentation methods: UNet [19], DeepLabv3+ [20], PraNet [22], CaraNet [24], TGANet [25], PVT-CASCADE [26], DuAT [28] and, SSFormer-L [29]. To ensure equitable comparison, identical hyperparameters were applied across all models. To evaluate the detection performance, we utilized established algorithms such as YOLOv10 [31], YOLOv9 [32], YOLOv8 [30], YOLOv7 [33], YOLOv6 [34], YOLOv5 [35]. Below, we present both segmentation and detection wise performance for modality and center-wise based data.
III-C Segmentation results on each modality of the dataset
III-C1 Results on BLI
In the BLI dataset, DuAT emerged as the top-performing model, achieving the highest mIoU of 0.6979 and mDSC of 0.8048. DuAT also excelled in recall with a leading score of 0.9082 and maintained a strong precision of 0.7647, resulting in the best F2 score of 0.8501. SSFormer-L followed closely with the second-highest mIoU of 0.6750, trailing by 2.29%. Both PVT-CASCADE and SSFormer-L provided close competition in mDSC, scoring 0.7873 and 0.7874, respectively. PraNet secured the second-best scores in recall (0.8876) and F2 (0.8348). Overall, DuAT demonstrated superior performance, with no other model consistently securing second-best across multiple metrics.
III-C2 Results on FICE
SSFormer-L dominated the FICE modality, achieving the highest mIoU of 0.7607 and mDSC of 0.8300, along with an impressive F2 score of 0.8526. Its recall score of 0.8713 was the second-best, while its precision score of 0.8013 remained competitive. PVT-CASCADE also performed well, with an mIoU of 0.7209 and mDSC of 0.7799. DuAT excelled in recall, achieving the highest score of 0.9082 for this modality, but its lower precision score of 0.5867 impacted its overall performance. Although DeepLabV3+ achieved the highest precision score of 0.9441, it did not lead in other metrics.
III-C3 Results on LCI
For the LCI dataset, SSFormer-L once again led the performance metrics, achieving an mIoU of 0.8567 and an mDSC of 0.9207. It attained a near-perfect precision score of 0.9466 and an impressive F2 score of 0.9106, making it the top choice for LCI segmentation. DuAT also performed exceptionally well, with an mIoU of 0.8551 and mDSC of 0.9194, leading in recall with a score of 0.9200 and delivering a strong precision score of 0.9247. PVT-CASCADE closely followed, showing balanced results across all metrics, particularly in recall (0.9074) and precision (0.9205). While TGANet exhibited high precision with a score of 0.9474, its slightly lower recall and mIoU scores prevented it from outperforming SSFormer-L and DuAT.
III-C4 Results on NBI
In the NBI dataset, segmentation models exhibited varying performance levels. PVT-CASCADE, based on PVTv2-B2, demonstrated superior performance with an mIoU of 0.7769, mDSC of 0.8586, and recall of 0.9385, highlighting its efficacy in polyp identification. Additionally, it achieved an F2 score of 0.8941, underscoring its dominance in this domain. SSFormer-L followed with the second-best performance, achieving an mIoU of 0.7608 and mDSC of 0.8432, alongside a strong recall of 0.9089, which was 2.96% lower than that of PVT-CASCADE. PraNet secured the highest precision score at 0.8836. The DuAT model also delivered competitive results, particularly notable in recall (0.8662) and precision (0.8741), although it did not surpass the comprehensive performance of PVT-CASCADE.
III-C5 Results on WLI
The WLI modality results were highly competitive, with SSFormer-L standing out as the top performer, achieving the best mIoU of 0.8821 and mDSC of 0.9294. SSFormer-L also led in recall with a score of 0.9314 and secured the second-best precision score of 0.9438, resulting in an impressive F2 score of 0.9288. PVT-CASCADE followed closely with an mIoU of 0.8731 and mDSC of 0.9219, demonstrating consistent performance with a recall of 0.9268 and precision of 0.9372. Although the performance gap between SSFormer-L and PVT-CASCADE was minimal, SSFormer-L’s slight edge in multiple metrics made it the best choice for WLI segmentation. The DuAT model also delivered strong results, with a mIoU of 0.8695 and mDSC of 0.9197, showcasing competitive recall and precision scores.
III-D Detection results on each modality of the dataset
III-D1 Results on BLI
In the BLI dataset, YOLOv9 stood out with the best MAP50, MAP50-95, and MAP75 scores of 0.688, 0.558, and 0.638, respectively. Precision was dominated by YOLOv8, which achieved a perfect score of 1, although it lagged behind in recall. YOLOv10 and YOLOv9 were tied for the best recall score of 0.5, indicating their strong performance in detecting positive cases.
III-D2 Results on FICE
@@
The FICE dataset results showed YOLOv10 outperforming other methods with the best MAP50, MAP50-95, and MAP75 scores of 0.887, 0.752, and 0.875, respectively. Additionally, YOLOv10 excelled in precision and recall, both achieving scores of 1 and 0.853, respectively, making it the most robust model for this modality.
III-D3 Results on LCI
@@
In the LCI dataset, YOLOv10 achieved the highest MAP50 and MAP75 scores of 0.995, although YOLOv9 closely followed with the best MAP50-95 score of 0.878. YOLOv10 also demonstrated superior performance in precision with a score of 1, while YOLOv9 achieved the best recall score of 1, highlighting its effectiveness in identifying true positive cases.
III-D4 Results on NBI
@@
For the NBI dataset, YOLOv9 delivered the best results with a MAP50 score of 0.688, a MAP50-95 score of 0.558, and a MAP75 score of 0.638. Precision was led by YOLOv8 with a perfect score of 1, but recall was highest for YOLOv10 and YOLOv9, both achieving a score of 0.5, showing their balanced performance in this modality.
III-D5 Results on WLI
@@
For the WLI dataset, YOLOv6 achieved the best MAP50 score of 0.925, while YOLOv8 and YOLOv5 shared the best MAP50-95 score of 0.766. The best MAP75 score of 0.868 was achieved by YOLOv8, which also demonstrated strong performance in precision with a score of 0.883. However, YOLOv8 excelled in recall, achieving the highest score of 0.88, closely followed by YOLOv5 and YOLOv7, both of which exhibited robust overall performance.
@@ IV Discussion
The quantitative results across the diverse datasets and modalities in PolypDB highlight the effectiveness of contemporary segmentation models, especially those utilizing advanced backbone architectures like PVTv2 and MiT-PLD. The variation in performance observed across modalities—NBI, WLI, BLI, FICE, and LCI—emphasizes the complex challenges inherent in polyp segmentation, where selecting the appropriate model architecture is crucial for attaining superior performance.

### v2

@@ III-A2 Implementation details
All the methods are implemented using the PyTorch 1.9 [18] framework, which is processed on an NVIDIA GeForce RTX 3090 system. We have used 80% of the dataset for training, 10% for validation, and the remaining 10% for testing. For polyp segmentation, we have first resized all the images into $512\times 512$ pixels. Next, we have used a minimal data augmentation, which includes random rotation, horizontal flipping, vertical flipping, and course dropout. A combination of binary cross-entropy loss and dice loss was selected as the loss function with the Adam optimizer, and a learning rate of $1e^{-4}$ was set. We have trained all the models with the same set of hyperparameters for 200 epochs (empirically set) with a batch size of 12. Early stopping and ReduceLROnPlateau was used to prevent the model from overfitting.
For polyp detection, we employed different hyperparameters tailored to optimize the performance of the detection algorithms. At first, we resized all the images to $640\times 480$ and used a simple data augmentation strategy which include random flipping, random rotation, random blur, mixup, mosaic and cutmix. All YOLO models were trained using a uniform set of hyperparameters, with the AdamW optimizer applied at a learning rate of $1\times 10^{-4}$ and a batch size of $16$. For federated segmentation experiments, all models are trained under the FedAvg algorithm [19] on the three centers. We use an AdamW optimizer [20] with a weight decay of 0.05 and momentum of 0.9. The models are trained with a mini-batch size of 32 for 100 epochs. The initial learning rate is 0.001, and the learning rate is reduced by a factor of 1/10 after each 30 epochs by using a cosine annealing schedule [21]. We normalize the images from each center using the mean and the standard variations of images from the same center.
@@ III-B Results
TABLE III: Comparison of quantitative results for segmentation on the PolypDB dataset. The best scores are shown in bold, whereas the second best score is underlined.
@@
Figure 2: Qualitative results for the detection task on the different modalities in the PolypDB dataset.
To evaluate the segmentation performance of the dataset, we employed several established segmentation methods.
III-B1 Segmentation results on each modality
Table VI shows the results of different segmentation methods on each modality of the dataset.
Results on BLI: In the BLI dataset, DuAT emerged as the top-performing model, achieving the highest mIoU of 0.6979 and mDSC of 0.8048. DuAT also demonstrated high recall with a score of 0.9082 and maintained a high precision of 0.7647, resulting in the best F2 score of 0.8501. SSFormer-L followed closely with the second-highest mIoU of 0.6750, trailing by 2.29%. Both PVT-CASCADE and SSFormer-L provided close competition in mDSC, scoring 0.7873 and 0.7848, respectively. PraNet secured the second-best scores in recall (0.8876) and F2 (0.8348). Overall, DuAT demonstrated superior performance, showcasing its segmentation capabilities across multiple metrics.
Results on FICE: SSFormer-L demonstrated the best results in the FICE modality, achieving the highest mIoU of 0.7607 and mDSC of 0.8300, along with an impressive F2 score of 0.8526. Its recall score of 0.8713 was the second-best, while its precision score of 0.8013 remained competitive. PVT-CASCADE also performed well, with an mIoU of 0.7209 and mDSC of 0.7799. DuAT excelled in recall, achieving the highest score of 0.9082 for this modality, but its lower precision score of 0.5867 impacted its overall performance. Although DeepLabV3+ achieved the highest precision score of 0.9441, it did not lead in other metrics.
Results on LCI: For the LCI dataset, SSFormer-L once again led the performance metrics, achieving an mIoU of 0.8567 and an mDSC of 0.9207. It attained a high precision score of 0.9466 and an impressive F2 score of 0.9106, making it the top choice for LCI segmentation. DuAT also performed exceptionally well, with an mIoU of 0.8551 and mDSC of 0.9194, leading in recall with a score of 0.9200 and delivering a strong precision score of 0.9247. PVT-CASCADE closely followed, showing balanced results across all metrics, particularly in recall (0.9074) and precision (0.9205). While TGANet exhibited a high precision of 0.9474, its slightly lower recall and mIoU scores prevented it from outperforming SSFormer-L and DuAT.
Results on NBI: In the NBI dataset, segmentation models exhibited varying performance levels. PVT-CASCADE, based on PVTv2-B2, demonstrated superior performance with an mIoU of 0.7769, mDSC of 0.8586, and recall of 0.9385, highlighting its efficacy in polyp identification. Additionally, it achieved an F2 score of 0.8941, underscoring its dominance in this domain. SSFormer-L followed with the second-best performance, achieving an mIoU of 0.7608 and mDSC of 0.8432, alongside a strong recall of 0.9089, which was 2.96% lower than that of PVT-CASCADE. PraNet secured the highest precision score at 0.8836. The DuAT model also delivered competitive results, particularly notable in recall (0.8662) and precision (0.8741), although it did not surpass the comprehensive performance of PVT-CASCADE.
Results on WLI: The WLI modality results were highly competitive, with SSFormer-L standing out as the top performer, achieving the best mIoU of 0.8821 and mDSC of 0.9294. SSFormer-L also led in recall with a score of 0.9314 and secured the second-best precision score of 0.9438, resulting in an impressive F2 score of 0.9288. PVT-CASCADE followed closely with an mIoU of 0.8731 and mDSC of 0.9219, demonstrating consistent performance with a recall of 0.9268 and precision of 0.9372. Although the performance gap between SSFormer-L and PVT-CASCADE was minimal, SSFormer-L’s slight edge in multiple metrics made it the best choice for WLI segmentation. The DuAT model also delivered strong results, with a mIoU of 0.8695 and mDSC of 0.9197, showcasing competitive recall and precision scores.
TABLE IV: Comparison of quantitative results for detection on the polyp datasets. The best scores are shown in bold, whereas the second best score is underlined.
Method mAP50 mAP50-95 mAP75 Precision Recall
@@ III-B1 Segmentation results on each modality
YOLOv8 [33] 0.659 0.502 0.559 1.000 0.318
YOLOv10 [34] 0.534 0.416 0.485 0.840 0.500
YOLOv9 [35] 0.688 0.558 0.638 0.846 0.500
@@
YOLOv10 [34] 0.887 0.752 0.875 1.000 0.853
YOLOv9 [35] 0.856 0.711 0.737 0.937 0.750
YOLOv7 [36] 0.734 0.642 0.734 0.856 0.750
@@
YOLOv8 [33] 0.833 0.771 0.833 1.000 0.667
YOLOv10 [34] 0.995 0.831 0.995 1.000 0.854
YOLOv9 [35] 0.972 0.878 0.972 0.857 1.000
@@
YOLOv8 [33] 0.659 0.502 0.559 1.000 0.318
YOLOv10 [34] 0.534 0.416 0.485 0.840 0.500
YOLOv9 [35] 0.688 0.558 0.638 0.846 0.500
@@
III-C Detection results on each modality of the dataset
Table IV illustrates the detection results on the PolypDB.
Results on BLI: In the BLI dataset, YOLOv9 achieves the highest performance with the best mAP50, mAP50-95, and mAP75 scores of 0.688, 0.558, and 0.638, respectively. In terms of precision, YOLOv8 surpasses other methods with a perfect score of 1.0000. However, for recall, both YOLOv9 and YOLOv10 achieve the same score of 0.5000. These results indicate strong performance in detecting positive cases, but the low recall scores highlight the challenge of improving the model’s ability to predict positive cases and reduce missed detections in this dataset.
@@ III-C Detection results on each modality of the dataset
Results on FICE: The FICE dataset results showed YOLOv10 outperforming other methods with the best mAP50, mAP50-95, and mAP75 scores of 0.8870, 0.7520, and 0.8750, respectively. Additionally, YOLOv10 excelled in precision and recall, achieving scores of 1.000 and 0.8530, respectively, making it the most robust model for this modality.
@@
Results on LCI: In the LCI dataset, YOLOv10 achieved the highest mAP50 and mAP75 scores of 0.9950, although YOLOv9 closely followed with the best mAP50-95 score of 0.8780. YOLOv10 also demonstrated superior performance in precision with a score of 1.000, while YOLOv9 achieved the best recall score of 1.0000, highlighting its effectiveness in identifying true positive cases.
@@
Results on NBI: For the NBI dataset, YOLOv9 delivered the best results with a mAP50 score of 0.6880, a mAP50-95 score of 0.5580, and a mAP75 score of 0.638. YOLOv8 achieves the highest precision score with 1.0000, but the YOLOv10 and YOLOv9 have the highest recall scores, both achieving a score of 0.5000, showing their balanced performance in this modality. In addition, based on the dataset characteristics, and the benchmark results, we can assume that this dataset is more challenging for the detection model to focus on the important feature of the polyp, thus leading to the redundant features learning (as shown in Figure 2), and low recall results.
@@
Results on WLI: For the WLI dataset, YOLOv6 achieved the highest mAP50 score of 0.9250, while YOLOv8 and YOLOv5 tied for the best mAP50-95 score of 0.7660. YOLOv8 also achieved the highest mAP75 score of 0.8680 and demonstrated strong precision with a score of 0.8830. Moreover, YOLOv8 excelled in the recall, achieving the top score of 0.8800, closely followed by YOLOv5 and YOLOv7. These results highlight the robustness of the dataset in guiding models to achieve high performance.
III-D Federated Segmentation Results on WLI
Table V presents the federated segmentation results. We apply the FedAvg algorithm [19] on three centers.
Average Results: On an average, SSFormer-L performed best with mIoU of 0.9214 and mDSC of 0.9550. It also achieved the highest precision and F2, with recall being the second best. SSFormer-L is followed closely by PVT-CASCADE and DuAT, which achieved very similar results. PVT-CASCADE excelled in the recall (0.9541), whereas DuAT reported the second-best mIoU and mDSC.
Results on BKAI: Similar to the average results, SSFormer-L attained the best outcomes on center zzz with mIoU of 0.9426 and mDSC of 0.9696. It also performed superior in precision and secured the second-highest scores in terms of recall and F2. PVT-CASCADE and DuAT closely matched their performance, where the former reported the highest recall (0.9757) and F2 (0.9713), and the latter achieved the second-highest mIoU and mDSC.
Results on Karolinska University Hospital, Stockholm, Sweden : On the center Karolinska University Hospital, Stockholm, Sweden data, DuAT was ranked as the top-performing model with the best scores in three metrics, including mIoU (0.9396), mDSC (0.9683) and F2 (0.9578). SSFormer-L proved to be the second-best performing model. Although DeepLabV3+ and TGANet excelled in precision and recall, respectively, their comparatively lower performance in other metrics prevented them from ranking among the best-performing models.
Results on Bærum Hospital, Vestre Viken Hospital Trust, Bærum, Norway: Showing consistently superior performance, SSFormer-L achieved the best outcomes in this case as well, with the highest mIoU, mDSC, precision and F2 of 0.9123, 0.9487, 0.9614, and 0.9439, respectively. The next best results are obtained using PVT-CASCADE, which are very similar to the DuAT outcomes.
TABLE V: Federated Segmentation results on WLI. Weights are aggregated equally. The best scores are shown in bold, whereas the second best score is underlined.
Method Backbone mIoU mDSC Recall Precision F2
Average
TransNetR [38] ResNet50 [24] 0.8771 0.9218 0.9289 0.9400 0.9236
U-Net [22] - 0.6362 0.7349 0.7929 0.7648 0.7562
U-NeXt [39] - 0.7049 0.7837 0.7977 0.8554 0.7853
DeepLabV3+ [23] ResNet50 [24] 0.9022 0.9423 0.9426 0.9539 0.9408
PraNet [25] Res2Net50 [26] 0.9048 0.9438 0.9390 0.9593 0.9397
CaraNet [27] Res2Net101 [26] 0.8941 0.9362 0.9324 0.9528 0.9325
TGANet [28] ResNet50 [24] 0.8837 0.9290 0.9445 0.9299 0.9345
PVT-CASCADE [29] PVTv2-B2 [30] 0.9105 0.9477 0.9541 0.9492 0.9504
DuAT [31] PVTv2-B2 [30] 0.9109 0.9478 0.9485 0.9550 0.9465
SSFormer-L [32] MiT-B4 [40] 0.9214 0.9550 0.9503 0.9642 0.9517
Center: BKAI
TransNetR [38] ResNet50 [24] 0.9199 0.9524 0.9598 0.9578 0.9561
U-Net [22] - 0.8072 0.8714 0.8788 0.9193 0.8729
U-NeXt [39] - 0.8018 0.8661 0.8660 0.9087 0.8635
DeepLabV3+ [23] ResNet50 [24] 0.9358 0.9656 0.9663 0.9678 0.9656
PraNet [25] Res2Net50 [26] 0.9331 0.9640 0.9669 0.9643 0.9654
CaraNet [27] Res2Net101 [26] 0.9198 0.9537 0.9613 0.9563 0.9578
TGANet [28] ResNet50 [24] 0.9311 0.9629 0.9608 0.9684 0.9612
PVT-CASCADE [29] PVTv2-B2 [30] 0.9369 0.9659 0.9757 0.9592 0.9713
DuAT [31] PVTv2-B2 [30] 0.9397 0.9679 0.9714 0.9661 0.9698
SSFormer-L [32] MiT-B4 [40] 0.9426 0.9696 0.9726 0.9685 0.9711
Center: Karolinska University Hospital, Stockholm, Sweden
TransNetR [38] ResNet50 [24] 0.8784 0.9308 0.8848 0.9899 0.9020
U-Net [22] - 0.6511 0.7287 0.6607 0.9820 0.6830
U-Next [39] - 0.6711 0.7414 0.7173 0.9449 0.7255
DeepLabV3+ [23] ResNet50 [24] 0.9159 0.9547 0.9225 0.9916 0.9349
PraNet [25] Res2Net50 [26] 0.8968 0.9428 0.9087 0.9847 0.9216
CaraNet [27] Res2Net101 [26] 0.8921 0.9406 0.9100 0.9802 0.9213
TGANet [28] ResNet50 [24] 0.8285 0.8959 0.9883 0.8340 0.9461
PVT-CASCADE [29] PVTv2-B2 [30] 0.9006 0.9438 0.9148 0.9829 0.9255
DuAT [31] PVTv2-B2 [30] 0.9396 0.9683 0.9510 0.9869 0.9578
SSFormer-L [32] MiT-B4 [40] 0.9186 0.9555 0.9257 0.9915 0.9370
Center: Bærum Hospital, Vestre Viken Hospital Trust, Bærum, Norway
TransNetR [38] ResNet50 [24] 0.8593 0.9091 0.9173 0.9314 0.9108
U-Net [22] - 0.5683 0.6798 0.7534 0.7074 0.7066
U-NeXt [39] - 0.6610 0.7465 0.7645 0.8342 0.7486
DeepLabV3+ [23] ResNet50 [24] 0.8876 0.9323 0.9326 0.9476 0.9302
PraNet [25] Res2Net50 [26] 0.8934 0.9356 0.9282 0.9569 0.9296
CaraNet [27] Res2Net101 [26] 0.8834 0.9289 0.9215 0.9499 0.9226
TGANet [28] ResNet50 [24] 0.8662 0.9161 0.9385 0.9153 0.9244
PVT-CASCADE [29] PVTv2-B2 [30] 0.8990 0.9399 0.9452 0.9442 0.9417
DuAT [31] PVTv2-B2 [30] 0.8985 0.9391 0.9399 0.9489 0.9372
SSFormer-L [32] MiT-B4 [40] 0.9123 0.9487 0.9416 0.9614 0.9439
Figure 3: Qualitative results for the different methods across various modalities in the PolypDB dataset.
@@ IV Discussion
The quantitative results across the diverse datasets and modalities in PolypDB highlight the effectiveness of contemporary segmentation models, especially those utilizing advanced backbone architectures like PVTv2 and MiT-B4. The variation in performance observed across modalities—NBI, WLI, BLI, FICE, and LCI—emphasizes the complex challenges inherent in polyp segmentation, where selecting the appropriate model architecture is crucial for attaining superior performance. Regarding polyp detection, the results from the YOLO family of models are promising, highlighting the quality of our dataset in supporting these models. Additionally, benchmark results and visualizations from our test sets reveal the challenges associated with boundary information learning in polyp detection. Specifically, the model’s performance on boundary detection and its ability to learn low-frequency features need to be improved to prevent it from missing small polyps or camouflaged objects where the patterns are similar to the background. We encourage further research to focus on solving this problem to improve polyp detection techniques.